OpenStudy (kirbykirby):

Stats/Probability Expectation: The sum of \(n\) independent squared standard normal random variables has a \(χ^2\) distribution with \(n\) degrees of freedom. Prove that if \(X\) has a \(χ^2\) distribution with \(n\) degrees of freedom, then \(E(X)=n\).

OpenStudy (kirbykirby):

This is what I did: Let \(X_i\)~\(N(0,1),i=1,2,...,n\) and let \(X=X_1^2+X_2^2+...+X_n^2\) So: \(E(X)=E(X_1^2+X_2^2+...+X_n^2)\) \(E(X)=E(X_1^2)+E(X_2^2)+...+E(X_n^2)\) Since each \(X_i^2\)~\(χ^2(1)\) <- 1 degree of freedom on chi-square, we know that \(E(X_i^2)=1\) since it's the expectation of a Chi-Square on 1 df. Hence, \(E(X)=1+1+...+1\) (n times 1) \(E(X)=n\) Someone told me though this might not be great because you are using the fact that the expectation of a Chi-Square on 1 df is 1, but we're trying to prove the case for df=n (but I dunno I didn't realize the Chi-Square's mean was it's degrees of freedom, just that in another class we used a Chi-Square on 1 df a lot and was told its mean was 1 as well...) Any other possible way to do it?